OpenStudy (anonymous):

differentiation question : show that the least value of x^2 + 2x(sinb) + 1 is cos^2b, where b is a constant.

OpenStudy (anonymous):

f'(x) = 2x + 2sin(b) = 0 for critical values x = -sinb plug this into f(x) = sin^2b -2sin^b + 1 = cos^2b

OpenStudy (anonymous):

simplify the eqn as (x+sinb)^2 + 1 - sinb^2 it will attain minimum value whwn the tyhing in the square is zero hence the minimum value is 1-sinb^b or cosb^2

OpenStudy (anonymous):

@jimmyrep my solution is better LOL

OpenStudy (anonymous):

yup - good solution